OpenStudy (anonymous):

What is 0.0085 as a percent

sam (.sam.):

0.85%

sam (.sam.):

100% = 1 So, 0.0085 = 0.0085 × 1 = 0.0085 × 100% = 0.85%.

Directrix (directrix):

0.0085 as a percent = .85 %

OpenStudy (anonymous):

.85%